# Trustworthy Retrosynthesis: Eliminating Hallucinations with a Diverse Ensemble of Reaction Scorers

### v2

@@ 4 Human Evaluation
We curated a dataset of over 4,500 reactions generated by our SSR models. Each reaction was evaluated and labeled by PhD-level chemists into one of the expert-defined categories, creating the first comprehensive dataset of its kind. This datased provides a way to evaluate the error patterns of our reaction plausibility scorers. A subset of this dataset will be released as a benchmark for the community.
@@ 5.1 Dataset
All of the models are based on the proprietary Pistachio (Mayfield et al. (2017)) dataset, either used as training data (generator, RP, RGP), or as a source of reference reactions (RRS). Pistachio offers substantial advantages over the commonly used USPTO-50K Schneider et al. (2016) and USPTO-FULL Dai et al. (2019b) datasets - it features enhanced curation, resulting in higher data quality and more comprehensive coverage of chemical reaction space. For training, we preprocess the dataset through a multi-step filtering pipeline that removes duplicate reactions, reactions from unrecognized reaction classes, entries with invalid SMILES, unmapped reactions, and reactions deemed unrelated to drug-like compound synthesis (too big molecules, "separation" reaction classes), retaining approximately 4 million reactions.
@@ 5.2 Path-Level Plausibility Evaluation
We compared the ability of each individual scorer (RP, RGP, and RRS) as well as the Meta-Scorer to filter out implausible reactions in multi-step retrosynthesis. We applied the scorers to a Transformer model trained on root-aligned reactions (Zhong et al., 2022) as a single-step generator (RootAligned).
The thresholds for filtering implausible reactions were selected based on results from the full chemist-evaluated reaction dataset described in 4, targeting 0.8 precision, which translated into approximately 90% of retrosynthetic pathways being successfully found. As baselines, we report results for RootAligned, for LocalRetro (Chen & Jung, 2021) without scoring, and for RootAligned with forward reaction scorer, enabling direct comparison with prior literature. For the search, we used the widely adopted Retro* algorithm based on the implementation from (Maziarz et al., 2025) with the expansion limit set to 500. For all systems, we used the same starting material database, eMolecules. Additionally, we compared against three publicly accessible retrosynthesis systems: IBM RXN (IBM, 2025), AiZynthFinder (Genheden et al., 2020b), and RetroChimera (Maziarz et al., 2024). AiZynthFinder was executed in its default configuration from the official repository (template generator + filtering based on a model trained on positive and artificial negative reactions), with the only modification being an increased time limit of 15 minutes to better match the runtime of other systems. IBM RXN was accessed through its free web application (IBM, 2025). RetroChimera, trained on Pistachio data and using the same eMolecules database, was queried via the Azure Foundry multi-step retrosynthesis endpoint (Microsoft, 2025). Due to a low per-call timeout, we repeated the queries multiple times for each target; thanks to prediction caching, this resulted in an effective 15-minute search time. It does not employ an explicit reaction scorer; plausibility is expected to be enforced implicitly by selecting the top-ranked reactions from the ensemble.
We evaluated the top-1 synthesis paths generated by all systems for 32 selected targets (listed in D). Each path was assigned a four-tier confidence score (Safe bet, Worthwhile, Rather not, Nonsense), determined by the lowest-scoring reaction in the path. This conservative scoring reflects the intuition that a single implausible step can invalidate an otherwise promising synthesis. Increasing the proportion of Safe bets, while eliminating Nonsense and reducing Rather Not paths is the goal of all retrosynthesis systems.
@@ 5.3 Reaction-Level Plausibility Prediction
Additionally, we compare the performance of individual scorers (RP, RGP and RRS) and the Meta-Scorer on individual reactions with ground truth labels established through expert chemist evaluations described in Section 4.
@@
$\text{overlap}=\frac{\left|\bigcap_{\text{scorer}\in\{\text{RGP},\text{RP},\text{RRS}\}}\text{FP}_{\text{scorer}}\right|}{\min_{\text{scorer}\in\{\text{RGP},\text{RP},\text{RRS}\}}\left|\text{FP}_{\text{scorer}}\right|},$ (3)
@@ 6.1 Path-Level Plausibility Evaluation
Pathway correctness comparison is presented in the figure 3. AiZynthFinder demonstrates significant limitations, failing to identify viable pathways for significant number of the target molecules while generating a substantial proportion of unreliable routes classified as Nonsense and Rather Not. IBM RXN shows improved performance by increasing the number of reliable pathways and reducing hallucinated predictions, yet fails to produce valid synthetic routes for a considerable fraction of target compounds. RetroChimera outputs pathways for the vast majority of targets. Although it produces a large percentage of Safe Bet pathways, it also generates a substantial number of Nonsense pathways. This shows that ensembling generators alone is not sufficient to mitigate errors for challenging targets.
Our baseline model (SSR generator without any filtering mechanism) significantly improves number of pathways found, providing solutions for all targets. However, confidence in its results is undermined by the significant presence of unreliable Nonsenense and Rather Not paths. Introducing individual scorers increases the fraction of targets for which no paths are found, a trade-off that can be desirable for the trustworthiness of the system — rejecting some targets is preferable to mixing reliable and unreliable pathways, as long as the remaining routes are correct. While RGP and RRS scorers reduce number of unreliable paths only modestly, our RP scorer demonstrates its value as a primary filter by eliminating all Nonsense reactions, though this comes at the cost of fewer Safe Bet and Worthwhile pathways. Finally, the Meta-Scorer delivers substantial improvements in reliability: significantly increasing Safe Bet paths, maintaining zero Nonsense results, and reducing Rather Not pathways. These results indicate that relying solely on individual scorers, especially those similar to commonly used approaches in the literature (such as RGP - feasibility classifier and forward scorers), is insufficient to achieve correct pathways. Only our RP-based solution combined with the Meta-Scorer provides routes with the highest overall plausibility while eliminating the most serious errors.
Figure 3: Comparison of our retrosynthesis systems against IBM RXN, AiZynthFinder, LocalRetro and RetroChimera. AiZynthFinder, IBM RXN, LocalRetro and RetroChimera fail on many targets. Our generator when used without any reaction scorer (indicated as "Our generator w/o scorer") finds pathways for all targets but includes unreliable routes. Introduction of individual scorers trades coverage for reliability, with RP eliminating all Nonsense pathways. The search system backed by the Meta-Scorer produces most trustworthy results.
@@ 6.2 Reaction-Level Plausibility Evaluation
Our results show that the Meta-Scorer outperforms individual scorers in both precision and recall, demonstrating effective integration of complementary signals. Figure 4 presents the ROC and precision-recall curves, with the Meta-Scorer achieving consistently higher area under the curve (AUC) values across both metrics. Similar curves broken down by reaction failure category can be found in Appendix E.
Figure 6 shows ROC-AUC values for each scorer broken down into different failure categories, illustrating that individual scorers demonstrate proficiency in filtering out reactions deemed implausible under different evaluation criteria. RGP achieves the best performance on Selectivity and Reactivity errors. RRS is most capable of detecting fundamental structural issues such as Reactant mismatches and Magic, in addition to One pot errors. RP shows a balanced profile, which explains its overall superior performance compared to RGP and RRS in Figure 4. By leveraging the unique strengths of each individual scorer, the Meta-Scorer maintains robust predictive performance across all failure categories.
We also analyze the overlap between false positives that each scorer fails to filter, as shown in Figure 6. The results show distinct complementarity: while RRS and RP exhibit high overlap in most categories, it is notably reduced for One-pot, Magic, and Reactant mismatch failure modes — the categories where Figure 6 demonstrates RRS’s superior performance. RGP and RP show consistently low overlap across all failure categories, indicating that these scorers capture different aspects of reaction implausibility. Importantly, when considering all three scorers jointly, the overlap drops to very low levels across all categories, providing strong evidence that each scorer contributes unique discriminative value essential for building a robust Meta-Scorer.
Figure 4: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the Meta-Scorer. The meta-scorer achieve higher AUC values for both ROC and PR curves, indicating better discrimination between plausible and implausible reactions. Among the individual scorers, RP shows the best performance.
@@ 7 Conclusions
In evaluating retrosynthesis systems and scorers, we made use of a novel labeling protocol where we leveraged chemists’ expertise to produce fine-grained labels for generated reactions. To our knowledge, this is the first instance of such a granular analysis of retrosynthetis systems’ output, where automated metrics and ad-hoc manual inspection were the norm. In order to facilitate further development in the field, we release the thirty-two targets used for path generation. While our evaluation process is generally applicable to retrosynthesis, RetroTrim was trained on data that biases it towards the medicinal chemistry context. Nonetheless, we hope that the insights and methodologies presented in this work lead to more reliable retrosynthesis in general.
@@ E.1 Magic
Figure 27: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the Meta-Scorer on Magic and No Problem reactions.
@@ E.2 Selectivity
Figure 28: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the Meta-Scorer on Selectivity and No Problem reactions.
@@ E.3 Functional group incompatibility
Figure 29: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the Meta-Scorer on Functional group incompatibility and No Problem reactions.
@@ E.4 Reactivity
Figure 30: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the Meta-Scorer on Reactivity and No Problem reactions.
@@ E.5 One pot
Figure 31: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the Meta-Scorer on One pot and No Problem reactions.
@@ E.6 Unstable
Figure 32: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the Meta-Scorer on Unstable and No Problem reactions.
@@ E.7 Reactants mismatch
Figure 33: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the Meta-Scorer on Reactants mismatch and No Problem reactions.
@@ Appendix F False Positives Counts By Failure Category
Figure 34: Counts of false positives produced by individual scorers versus the Meta-Scorer across different failure categories, with sample sizes indicated for each category.
@@ Appendix G True Negatives Counts By Failure Category
Figure 35: Counts of true negatives produced by individual scorers versus the Meta-Scorer across different failure categories, with sample sizes indicated for each category.

### v3

@@ 4 Human Evaluation
We curated a dataset of over 4,500 reactions generated by our SSR models. Each reaction was evaluated and labeled by PhD-level chemists into one of the expert-defined categories, creating the first comprehensive dataset of its kind. This datased provides a way to evaluate the error patterns of our reaction plausibility scorers. A subset of 500 reactions from this dataset will be released as a benchmark for the community.
@@ 5.1 Dataset
All of our generators and scorers are based on the proprietary Pistachio (2024Q3 release) (mayfield2017pistachio) dataset, either used as training data (RootAligned, RP, RGP), or as a source of reference reactions (RRS). Pistachio offers substantial advantages over the commonly used USPTO-50K (uspto) and USPTO-FULL (uspto_full) datasets - it features enhanced curation, resulting in higher data quality and more comprehensive coverage of chemical reaction space. For training, we preprocess the dataset through a multi-step filtering pipeline that removes duplicate reactions, reactions from unrecognized reaction classes, entries with invalid SMILES, unmapped reactions, and reactions deemed unrelated to drug-like compound synthesis (molecules with >100 atoms, "separation" reaction classes), retaining approximately 4 million reactions.
@@ 5.2 Path-Level Plausibility Evaluation
We compared the ability of each individual scorer (RP, RGP, and RRS) as well as the MetaScorer at filtering out implausible reactions in multi-step retrosynthesis. All scorer variants were used together with our BART-based RootAligned generator.
The thresholds for filtering implausible reactions were maximized under the constraint that the resulting system finds paths for at least  90% of targets. In practice, this corresponded to a precision value of  0.8 on the reaction dataset described in 4. We targeted both a relatively strict threshold (to keep precision high) and broad coverage in terms of targets for which routes are found. Importantly, these thresholds were not tuned to optimize benchmark performance. As baselines, we report results for RootAligned without scoring, RootAligned with a forward reaction scorer, and LocalRetro (LocalRetro). The forward scorer is implemented in the same BART architecture as RootAligned, except it is trained to predict products based on substrates. For the search, we used the widely adopted Retro* algorithm based on the implementation from (MaziarzSyntheseus) with the expansion limit set to 500. For all systems, we used the same starting material database, eMolecules.
Additionally, we compared against three publicly accessible retrosynthesis systems: AiZynthFinder (genheden2020aizynthfinder), IBM RXN (ibmrxnapp), and RetroChimera (maziarz2024chimera). AiZynthFinder was used in its default configuration from the official repository, which includes a template generator with a filtering model trained to distinguish valid reactions from artificially generated negatives. The only modification we make is an increased time limit of 15 minutes to better match the runtime of other systems. IBM RXN, which makes use of a forward model in its search, was accessed through its free web application (ibmrxnapp). RetroChimera was queried via the Azure Foundry multi-step retrosynthesis endpoint (RetroChimeraApp). Due to a low per-call timeout, we repeated the queries multiple times for each target; thanks to prediction caching, this resulted in an effective 15-minute search time. Like RetroTrim, RetroChimera does not employ an explicit reaction scorer, instead it aims to enforce plausibility by selecting the top-ranked reactions from an ensemble of generators.
We evaluated the top-1 synthesis paths generated by all systems for 32 selected targets (listed in D). Each reaction in the generated paths was manually evaluated by expert chemists according to the evaluation protocol described in 4. Each path was assigned a four-tier confidence score (Safe bet, Worthwhile, Rather not, Nonsense), determined by the lowest-scoring reaction in the path. This conservative scoring reflects the intuition that a single implausible step can invalidate an otherwise promising synthesis. Increasing the proportion of Safe bets, while eliminating Nonsense and reducing Rather Not paths is the goal of all retrosynthesis systems.
@@ 5.3 Reaction-Level Plausibility Prediction
Additionally, we compare the performance of individual scorers (RP, RGP and RRS) and the MetaScorer on individual reactions with ground truth labels established through expert chemist evaluations described in Section 4.
@@
$\text{overlap}=\frac{\left|\bigcap_{\texttt{scorer}\in\{\texttt{RGP},\texttt{RP},\texttt{RRS}\}}\text{FP}_{\texttt{scorer}}\right|}{\min_{\texttt{scorer}\in\{\texttt{RGP},\texttt{RP},\texttt{RRS}\}}\left|\text{FP}_{\texttt{scorer}}\right|},$ (3)
@@ 6.1 Path-Level Plausibility Evaluation
Pathway correctness comparison is presented in the figure 3. AiZynthFinder demonstrates significant limitations, failing to identify viable pathways for significant number of the target molecules while generating a substantial proportion of unreliable routes classified as Nonsense and Rather Not. IBM RXN shows improved performance by increasing the number of reliable pathways and reducing hallucinated predictions, yet fails to produce valid synthetic routes for a considerable fraction of target compounds. RetroChimera outputs pathways for the vast majority of targets. Although it produces a large percentage of Safe Bet pathways, it also generates a substantial number of Nonsense pathways, showing that ensembling current generators alone is not sufficient to mitigate errors for challenging targets.
RetroAligned without scoring significantly improves number of pathways found, providing solutions for all targets. However, confidence in its results is undermined by the significant presence of unreliable Nonsenense and Rather Not paths. Introducing individual scorers increases the fraction of targets for which no paths are found, a trade-off that can be desirable for the trustworthiness of the system — rejecting some targets is preferable to mixing reliable and unreliable pathways, as long as the remaining routes are correct. While RGP and RRS scorers reduce number of unreliable paths only modestly, our RP scorer demonstrates its value as a primary filter by eliminating all Nonsense reactions, though this comes at the cost of fewer Safe Bet and Worthwhile pathways. Finally, RetroTrim, utilizing the MetaScorer, delivers substantial improvements in reliability: significantly increasing Safe Bet paths, maintaining zero Nonsense results, and reducing Rather Not pathways. In our results, individual scorers similar to those commonly appearing in the literature (such as RGP - feasibility classifier and forward scorers), weren’t sufficient to achieve correct pathways. RetroTrim provides the largest number of problematic routes while eliminating the most serious errors.
Figure 3: Comparison of our retrosynthesis generator (RootAligned) with different scorers against IBM RXN, AiZynthFinder, LocalRetro, and RetroChimera. Among AiZynthFinder, IBM RXN, LocalRetro and RetroChimera, RetroChimera performs significantly better than others, but it still fails on >25% targets, with a significant number of hallucinations. RootAligned without any reaction scorer finds pathways for all targets but includes unreliable routes. Introduction of individual scorers trades coverage for reliability, with RP eliminating all Nonsense pathways. RetroTrim, backed by the MetaScorer produces the most trustworthy results.
@@ 6.2 Reaction-Level Plausibility Evaluation
Our results show that the MetaScorer outperforms individual scorers in both precision and recall, demonstrating effective integration of complementary signals. Figure 4 presents the ROC and precision-recall curves, with the MetaScorer achieving consistently higher area under the curve (AUC) values across both metrics. Similar curves broken down by reaction failure category can be found in Appendix E.
Figure 4: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the MetaScorer. The MetaScorer achieve higher AUC values for both ROC and PR curves, indicating better discrimination between plausible and implausible reactions. Among the individual scorers, RP shows the best performance.
Figure 6 shows ROC-AUC values for each scorer broken down into different failure categories, illustrating that individual scorers demonstrate proficiency in filtering out reactions deemed implausible under different evaluation criteria. RGP achieves the best performance on Selectivity and Reactivity errors. RRS is most capable of detecting fundamental structural issues such as Reactant mismatches and Magic, in addition to One pot errors. RP shows a balanced profile, which explains its overall superior performance compared to RGP and RRS in Figure 4. By leveraging the unique strengths of each individual scorer, the MetaScorer maintains robust predictive performance across all failure categories.
@@
We also analyze the overlap between false positives that each scorer fails to filter, as shown in Figure 6. The results show distinct complementarity: while RRS and RP exhibit high overlap in most categories, it is notably reduced for One-pot, Magic, and Reactant mismatch failure modes — the categories where Figure 6 demonstrates RRS’s superior performance. RGP and RP show consistently low overlap across all failure categories, indicating that these scorers capture different aspects of reaction implausibility. Importantly, when considering all three scorers jointly, the overlap drops to very low levels across all categories, providing strong evidence that each scorer contributes unique discriminative value essential for building a robust MetaScorer.
@@ 7 Conclusions
In evaluating retrosynthesis systems and scorers, we made use of a novel labeling protocol where we leveraged chemists’ expertise to produce fine-grained labels for generated reactions. To our knowledge, this is the first instance of such a granular analysis of retrosynthetis systems’ output, where automated metrics and ad-hoc manual inspection were the norm. In order to facilitate further development in the field, we release the thirty-two targets used for path generation. While our evaluation process is generally applicable to retrosynthesis, RetroTrim was trained on data that biases it towards the medicinal chemistry context. We also note that by focusing on the top-1 performance of retrosynthesis systems, we leave open for further work the analysis of how different plausibility filtering methods impact the diversity of resulting paths. Nonetheless, we hope that the insights and methodologies presented in this work lead to more reliable retrosynthesis in general.
@@ E.1 Magic
Figure 27: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the MetaScorer on Magic and No Problem reactions.
@@ E.2 Selectivity
Figure 28: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the MetaScorer on Selectivity and No Problem reactions.
@@ E.3 Functional group incompatibility
Figure 29: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the MetaScorer on Functional group incompatibility and No Problem reactions.
@@ E.4 Reactivity
Figure 30: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the MetaScorer on Reactivity and No Problem reactions.
@@ E.5 One pot
Figure 31: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the MetaScorer on One pot and No Problem reactions.
@@ E.6 Unstable
Figure 32: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the MetaScorer on Unstable and No Problem reactions.
@@ E.7 Reactants mismatch
Figure 33: ROC (on the left) and precision-recall (on the right) curves comparing the performance of individual scorers versus the MetaScorer on Reactants mismatch and No Problem reactions.
@@ Appendix F False Positives Counts By Failure Category
Figure 34: Counts of false positives produced by individual scorers versus the MetaScorer across different failure categories, with sample sizes indicated for each category.
@@ Appendix G True Negatives Counts By Failure Category
Figure 35: Counts of true negatives produced by individual scorers versus the MetaScorer across different failure categories, with sample sizes indicated for each category.